One of the reactions used to inflate automobile air bags involves sodium azide



(NaN3). The balanced equation will be: 2NaN3(s) → 2Na(s) + 3N2(g). Determine the mass of N2



produced from the decomposition of 100.0 g of NaN3.

1
Expert's answer
2022-05-20T10:46:03-0400

The mol number of sodium azide:

"n(NaN_3) = \frac{m(NaN_3)}{M(NaN_3)}=\frac{100.0 g}{65.0g } = 1.54 mol"

By the balanced reaction we observed following:

"2n(NaN_3) = 3n(N_2)"

"n(N_2) = \frac{2}{3}n(NaN_3)"

"m(N_2) = n(N_2) \times M(N_2) = \frac{2}{3}n(NaN_3) \times M(N_2)"

"m(N_2) = \frac{2}{3} \times 1.54 mol \times 14 g/mol = 14.4 g"